### v1

@@ 6 Conditions for Coordinatewise Polarization
Let $P$ be a full support prior over $\Theta$ and $Q$ the posterior after receiving a set signal with realization $\Gamma$. Let $P|_{\Gamma},{P}|_{\Gamma^{C}}$ be the conditional distributions of $P$ over $\Gamma$ and $\Gamma^{C}$ respectively. Then $Q\preceq_{cw}P\iff P|_{\Gamma}\preceq_{cw}{P}|_{\Gamma^{C}}$.
If an identified set violates one of the first two conditions of Theorem 5 then either the low agent’s beliefs cannot move down in the coordinatewise order or the high agent’s beliefs cannot move up. For example, if $\Gamma$ is biased downward, then it cannot be the case that $P^{H}\prec_{cw}Q^{H}$. If an identified set satisfies the first two conditions but violates the third then, while it is possible that the low agent’s beliefs move down, and possible that the high agent’s beliefs move up, it cannot be the case that both occur for the same identified set.
To illustrate, consider again the $2\times 2$ state space. Here the only non-compensatory set that satisfies the first two conditions is the “off-diagonal” $\Gamma=\{(x_{1},y_{2}),(x_{2},y_{1})\}$. Assume towards contradiction that $\Gamma$ can polarize. By Lemma 6 and the fact that the coordinatewise order is generated by sums of univariate increasing functions, for any such function $u$ it must be the case that $E^{L}[u|\Gamma]\leq E^{L}[u|\Gamma^{C}]$ and $E^{H}[u|\Gamma]\geq E^{H}[u|\Gamma^{C}]$. Consider the function $u(\theta)=u_{1}(\theta_{1})+u_{2}(\theta_{2})$, where $u_{1}(x_{1})=u_{2}(y_{1})=0$, and $u_{1}(x_{2})=u_{2}(y_{2})=1$. The values of $u$ over all states in $\Theta$ are summarized in Figure 6. Since $u(\theta)=1$ for all $\theta\in\Gamma$, we get $E^{L}[u|\Gamma]=E^{H}[u|\Gamma]=1$. So the above implies that $E^{L}[u|\Gamma^{C}]\geq 1$ and $E^{H}[u|\Gamma^{C}]\leq 1$. But then the unconditional expectations satisfy $E^{L}[u]\geq 1\geq E^{H}[u]$. However, the assumption $P^{L}\prec^{*}_{cw}P^{H}$ implies $E^{L}[u]<E^{H}[u]$, a contradiction. [^19] The final inequality follows from Lemma 11 in the appendix. The reason we need to assume that $P^{H}$ strongly dominates $P^{L}$ is to ensure that this inequality is strict.
[…]
Figure 6: Values of $u(\theta)$. States in the off-diagonal $\Gamma$ are outlined in red.
The proof that the non-compensatory condition is necessary is a straightforward extension of this example and is relegated to the appendix along with the rest of the necessity direction. The sufficiency proof is more involved. Its outline is presented here, with intermediate lemmas proved in the appendix. The key step is Lemma 6, which is proved by induction on the size of the state space. Although the sufficiency proof does not provide the polarizing priors explicitly, the induction argument of Lemma 6 can be adapted to construct them recursively.
Proof of Theorem 5 (sufficiency). Since $\Theta\subset\mathbb{R}^{2}$ we can assume $\Theta=\{x_{1},\ldots,x_{m}\}\times\{y_{1},\ldots,y_{n}\}$ for $m,n\geq 2$. Throughout this proof, let $F^{L}_{i},F^{H}_{i}$ denote the cdfs of the marginal distributions of $P^{L}$ and $P^{H}$, respectively on the $i$ coordinate. Let $F^{L}_{i}|_{\Gamma},F^{L}_{i}|_{\Gamma}^{C}$ denote the cdfs of the marginal distributions of $P^{L}|_{\Gamma}$ and $P^{L}|_{\Gamma^{C}}$, and similarly define $F^{H}_{i}|_{\Gamma},F^{H}_{i}|_{\Gamma^{C}}$.
@@
Let $\Gamma\subset\Theta$ be a set that satisfies all three conditions. We will establish the existence of $P^{L},P^{H}\in\Delta(\Theta)$ such that strong coordinatewise polarization occurs upon observing $\Gamma$. The proof proceeds in three steps.
@@
Recall that a subset of a partially ordered set is an antichain if any two distinct elements in the set are incomparable to each other. We will consider antichains that are subsets of $\Theta$, under the order $\leq$ defined above. We first define a partial order over such antichains.
@@
By Lemma 6, to verify that such distributions exist, it suffices to prove the following.
@@
For any set $A\subset\Theta$, let $U_{A}$ denote the uniform distribution over $A$, and let $F^{U}_{i}$ denote its marginal cdf on the $i$ coordinate. Let $\delta\in(0,1)$ and define [^21] These distributions are defined as weighted mixtures of their summands. For example, for every $\theta\in\Theta$: $P^{L}(\theta)=(1-\delta)\left[(1-\delta)P^{L}|_{\min(\Gamma)}(\theta)+\delta U_{\Gamma\setminus\min(\Gamma)}(\theta)\right]+\delta\left[(1-\delta)P^{L}|_{\max(\Gamma^{C})}(\theta)+\delta U_{\Gamma^{C}\setminus\max(\Gamma^{C})}(\theta)\right].$
@@
With these definitions of the priors, the following lemma establishes polarization, completing the sufficiency proof.
Lemma 8.
Polarization occurs for the identified set $\Gamma$ and the priors in (2).
7 Final Remarks
This paper focused on the possibility of belief polarization among Bayesian agents and on conditions for it to arise. We conclude by briefly discussing two additional issues: the probability of polarization, and the relationship between polarization of beliefs and polarization of actions.
7.1 Probability of Polarization

### v2

@@ A.1 Proof of Theorem 2
The conditions are necessary:
  1. (i)
    $\Gamma$ cannot polarize if it is not spanning or if $\Gamma^{C}$ is not spanning: Assume that $\Gamma$ is not spanning and assume towards contradiction that coordinatewise polarization occurs. If $\Gamma$ does not obtain the minimal value on one of the coordinates, w.l.o.g the first, then $F^{L}_{1}|_{\Gamma}(x_{1})=0<F^{L}_{1}|_{\Gamma^{C}}(x_{1})$. By Lemma 2, this violates $Q^{L}\prec_{cw}P^{L}$, a contradiction. Similarly, if $\Gamma$ does not obtain the maximal value on the first coordinate then $F^{H}_{1}|_{\Gamma}(x_{m-1})=1<F^{H}_{1}|_{\Gamma^{C}}(x_{m-1})$, a contradiction to $P^{H}\prec_{cw}Q^{H}$. Similar arguments show that $\Gamma^{C}$ must be spanning as well.
  2. (ii)
    Note that $u(\theta)=0$ for all $\theta\leq\theta^{*}$, $u(\theta)=2$ for all $\theta\gg\theta^{*}$, and $u(\theta)=1$ otherwise. So, the assumption that $\Gamma$ does not contain any state weakly below $\theta^{*}$ and contains all states strictly above it implies that $E^{L}[u|\Gamma]>1$ and $E^{L}[u|\Gamma^{C}]<1$. [^26] Recall that these are the conditional expectations according to the prior $P^{L}$. Since $\theta^{*}\ll\bar{\theta}$, there exist states $\theta\in\Theta$ with $\theta\gg\theta^{*}$. So, the assumption that $P^{L}$ has full support implies the strict inequalities. Since $u$ is a sum of univariate increasing functions, and the coordinatewise order is generated by such functions (Lemma 4), this violates $P^{L}|_{\Gamma}\prec_{cw}P^{L}|_{\Gamma^{C}}$, and therefore by Lemma 2 violates $Q^{L}\prec_{cw}P^{L}$, a contradiction.
  3. (iii)
    $\Gamma$ cannot polarize if it is compensatory: Assume that $\Gamma$ is compensatory and assume towards contradiction that coordinatewise polarization occurs. Let $\theta^{*}\in\Theta$ be such that $\theta^{*}\ll\bar{\theta}$ and such that for every $\theta^{\prime}\in\Gamma$ it is not the case that $\theta^{\prime}\leq\theta^{*}$ and it is not the case that $\theta^{\prime}\gg\theta^{*}$. Then, for the function $u$ defined in the previous step we have $u(\theta)\equiv 1$ for every $\theta\in\Gamma$. Next, by Lemma 7, the strong coordinatewise order implies strict inequality of expectations for nontrivial sums of univariate increasing functions. That is, $P^{L}\prec^{*}_{cw}P^{H}$ and the fact that $u$ is not constant over $\Theta$ imply that $E^{L}[u]<E^{H}[u]$. By Lemma 2, $Q^{L}\prec_{cw}P^{L}$ implies $P^{L}|_{\Gamma}\prec_{cw}P^{L}|_{\Gamma^{C}}$, which implies
    $1=E^{L}[u|\Gamma]\leq E^{L}[u|\Gamma^{C}]$
    Therefore,
    Similarly, $P^{H}\prec_{cw}Q^{H}$ implies
    $1=E^{H}[u|\Gamma]\geq E^{H}[u|\Gamma^{C}]$
@@
Let $\Gamma\subset\Theta$ be a set that satisfies all three conditions. We will establish the existence of priors $P^{L},P^{H}\in\Delta(\Theta)$ such that strong coordinatewise polarization occurs upon observing $\Gamma$. The proof proceeds in three steps.
@@
Recall that a subset of a partially ordered set is an antichain if any two distinct elements in the set are incomparable to each other. We will consider antichains that are subsets of $\Theta$, under the order $\leq$ defined above. We first define a binary dominance relation over such antichains.
@@
Remark 3.
Our original definition of the order $\succeq^{*}_{cw}$ applied to two distributions fully supported on $\Theta$. Here, the distributions’ supports are strictly contained in $\Theta$ but we still take $\succeq^{*}_{cw}$ to imply the strict marginal inequalities over $\Theta$. That is, $P_{A}\prec^{*}_{cw}P_{B}$ means that $F^{A}_{1}(x)>F^{B}_{1}(x)$ for each $x\in\Theta_{1}$ with $x<x_{m}$ and $F^{A}_{2}(y)>F^{B}_{2}(y)$ for each $y\in\Theta_{2}$ with $y<y_{n}$.
@@
We will make use of the following result.
@@
We show that we can find distributions that satisfy all the conditions in (2) by considering four cases depending on whether or not $\Gamma$ contains the extreme states $\underaccent{\bar}{\theta},\bar{\theta}$.
Case 1: $\underaccent{\bar}{\theta},\bar{\theta}\in\Gamma$. In this case, define $P^{L}|_{\min(\Gamma)}$ and $P^{H}|_{\max(\Gamma)}$ as Dirac measures on the singleton on which they are supported: $P^{L}|_{\min(\Gamma)}=D_{\underaccent{\bar}{\theta}}$ and $P^{H}|_{\max(\Gamma)}=D_{\bar{\theta}}$. It is immediate that $P^{L}|_{\min(\Gamma)}\prec^{*}_{cw}P^{H}|_{\max(\Gamma)}$.
Define $P^{L}|_{\max(\Gamma^{C})}$ and $P^{H}|_{\min(\Gamma^{C})}$ as uniform distributions on their respective supports. By the complement spanning condition, $\max(\Gamma^{C})$ obtains the maximal value on each coordinate, which implies that for all $i\in[m-1]$ and $j\in[n-1]$,
$\begin{split}F^{L}_{1}|_{\min(\Gamma)}(x_{i})&=1>F^{L}_{1}|_{\max(\Gamma^{C})}(x_{i})\\ F^{L}_{2}|_{\min(\Gamma)}(y_{j})&=1>F^{L}_{2}|_{\max(\Gamma^{C})}(y_{j})\end{split}$
Hence, $P^{L}|_{\min(\Gamma)}\prec^{*}_{cw}P^{L}|_{\max(\Gamma^{C})}$. A similar argument shows $P^{H}|_{\min(\Gamma^{C})}\prec^{*}_{cw}P^{H}|_{\max(\Gamma)}$, so the distributions satisfy (2). [^28] The resulting priors $P^{L},P^{H}$ in this case are going to be those constructed in the proof of Theorem 1.
Case 2: $\underaccent{\bar}{\theta},\bar{\theta}\in\Gamma^{C}$. Define $P^{H}|_{\min(\Gamma^{C})}=D_{\underaccent{\bar}{\theta}}$ and $P^{L}|_{\max(\Gamma^{C})}=D_{\bar{\theta}}$. By Lemma 9 we have $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma)$. So, by Lemma 8, there exist $P^{L}|_{\min(\Gamma)}$ fully supported on $\min(\Gamma)$ and $P^{H}|_{\max(\Gamma)}$ fully supported on $\max(\Gamma)$ such that $P^{L}|_{\min(\Gamma)}\prec^{*}_{cw}P^{H}|_{\max(\Gamma)}$. Similar to Case 1, the spanning condition together with the fact that $P^{L}|_{\min(\Gamma)},P^{H}|_{\max(\Gamma)}$ are fully supported implies that $P^{L}|_{\min(\Gamma)}\prec^{*}_{cw}P^{L}|_{\max(\Gamma^{C})}$ and $P^{H}|_{\min(\Gamma^{C})}\prec^{*}_{cw}P^{H}|_{\max(\Gamma)}$ so (2) holds.
Case 3: $\underaccent{\bar}{\theta}\in\Gamma,\bar{\theta}\in\Gamma^{C}$. Define $P^{L}|_{\min(\Gamma)}=D_{\underaccent{\bar}{\theta}}$ and $P^{L}|_{\max(\Gamma^{C})}=D_{\bar{\theta}}$. It is immediate that $P^{L}|_{\min(\Gamma)}\prec^{*}_{cw}P^{L}|_{\max(\Gamma^{C})}$. By Lemma 9 we have $\min(\Gamma^{C})<^{\Theta}_{AC}\max(\Gamma)$. So, by Lemma 8, there exist $P^{H}|_{\min(\Gamma^{C})}$ and $P^{H}|_{\max(\Gamma)}$, each with full support, such that $P^{H}|_{\min(\Gamma^{C})}\prec^{*}_{cw}P^{H}|_{\max(\Gamma)}$. Similar to Case 1, the spanning condition and the fact that $P^{H}|_{\max(\Gamma)}$ has full support implies that $P^{L}|_{\min(\Gamma)}\prec^{*}_{cw}P^{H}|_{\max(\Gamma)}$, so (2) holds.
Case 4: Follows from Case 3 by swapping $\Gamma\leftrightarrow\Gamma^{C}$ and $L\leftrightarrow H$.
@@
For any set $A\subset\Theta$, let $U_{A}$ denote the uniform distribution over $A$. Let $\delta\in(0,1)$ and define [^29] These distributions are defined as weighted mixtures of their summands. The sets $\Gamma^{C}\setminus\max(\Gamma^{C})$ and $\Gamma^{C}\setminus\min(\Gamma^{C})$ may be empty. This occurs when $\Gamma^{C}=\min(\Gamma^{C})=\max(\Gamma^{C})$. In this case, define the corresponding prior as a mixture over the remaining sets, i.e., $P^{L}=(1-\delta)\left[(1-\delta)P^{L}|_{\min(\Gamma)}+\delta U_{\Gamma\setminus\min(\Gamma)}\right]+\delta P^{L}|_{\max(\Gamma^{C})}$ and similarly for $P^{H}$. Both priors are fully supported on $\Theta$ and it is straightforward to verify that all inequalities below hold for this case. The remaining sets in (3) cannot be empty: the min and max sets of any finite set are nonempty, and if it were the case that $\Gamma=\min(\Gamma)=\max(\Gamma)$, then all states in $\Gamma$ would be incomparable to each other, which implies that $\Gamma$ is compensatory.
@@
These priors are fully supported on $\Theta$ by construction. We will show that for sufficiently small $\delta$, polarization occurs for the identified set $\Gamma$ and the priors in (3).
For any $A\subset\Theta$ let $F^{U}_{i}|_{A}$ denote the cdf of the marginal of $U_{A}$ on the $i$ coordinate. Note that $P^{L}|_{\Gamma}=(1-\delta)P^{L}|_{\min(\Gamma)}+\delta U_{\Gamma\setminus\min(\Gamma)}$ and $P^{L}|_{\Gamma^{C}}=(1-\delta)P^{L}|_{\max(\Gamma^{C})}+\delta U_{\Gamma^{C}\setminus\max(\Gamma^{C})}$. Thus, for every $i\in[m-1]$ and for $\delta$ sufficiently small such that $(1-\delta)\varepsilon>\delta$ we have
Similarly, $F^{L}_{2}|_{\Gamma}(y_{j})>F^{L}_{2}|_{\Gamma^{C}}(y_{j})$ for every $j\in[n-1]$. So $P^{L}|_{\Gamma}\prec_{cw}P^{L}|_{\Gamma^{C}}$. An analogous argument shows that $P^{H}|_{\Gamma}\succ_{cw}P^{H}|_{\Gamma^{C}}$. Thus, by Lemma 2, for sufficiently small $\delta$ we have $Q^{L}\prec_{cw}P^{L}$ and $P^{H}\prec_{cw}Q^{H}$.
So,
Proof of Lemma 6. Let $F^{L}$ and $F^{H}$ be the corresponding cdfs. Denote $X=\{x_{1},\ldots,x_{n}\}$, where $x_{1}<\ldots<x_{n}$. Using summation by parts,
Since $P^{L}\prec^{*}_{st}P^{H}$, we have $F^{L}(x_{i})-F^{H}(x_{i})>0$ for every $i\in[n-1]$. This, together with the fact that $u$ is weakly increasing and not constant implies that the summands in the final expression are weakly greater than 0 for all $i\in[n-1]$ and strictly greater than 0 for some $i\in[n-1]$, hence $E^{H}[u]-E^{L}[u]>0$.
Proof of Lemma 7. Since $u$ is not constant, one of the functions $u_{i}$ must not be constant. Thus, since $P^{L}_{i}\prec^{*}_{st}P^{H}_{i}$ for all $i\in[d]$, Lemma 6 implies that $E_{P^{L}_{i}}[u_{i}]\leq E_{P^{H}_{i}}[u_{i}]$ for all $i\in[d]$ and $E_{P^{L}_{j}}[u_{j}]<E_{P^{H}_{j}}[u_{j}]$ for at least one $j\in[d]$. So,
$E^{L}[u]=\sum_{i=1}^{d}E_{P^{L}_{i}}[u_{i}]<\sum_{i=1}^{d}E_{P^{H}_{i}}[u_{i}]=E^{H}[u]$
Base case: Assume $\Theta=\{x_{1},x_{2}\}\times\{y_{1},y_{2}\}$. Let $A,B\subset\Theta$ be two nonempty antichains such that $B>^{\Theta}_{AC}A$. Note that, apart from the two-element antichain $\{(x_{2},y_{1}),(x_{1},y_{2})\}$, all other antichains in $\Theta$ are singletons. Hence, at least one of the disjoint antichains $A,B$ is a singleton. Assume w.l.o.g. that this is $B$. Since $B$ hits the maximal value on both coordinates, it must be the case that $B=\{(x_{2},y_{2})\}$. Let $P^{B}$ be the Dirac measure on $B$ and $P^{A}$ be any full support measure on $A$. Then $F^{A}_{1}(x_{1})>0=F^{B}_{1}(x_{1})$ and $F^{A}_{2}(y_{1})>0=F^{B}_{2}(y_{1})$ so $P^{A}\prec^{*}_{cw}P^{B}$.
Inductive step: Assume the induction hypothesis holds for all composite numbers $s^{\prime}<s$ and consider a state space $\Theta$ with $|\Theta|=s$. Let $A,B\subset\Theta$ be two nonempty antichains such that $B>^{\Theta}_{AC}A$. If one of these is a singleton then the proof is straightforward by an argument similar to the base case. Assume that $|A|,|B|\geq 2$. We can write $A=\{\delta^{1},\ldots,\delta^{k}\},B=\{\theta^{1},\ldots,\theta^{\ell}\}$, where both sets are ordered according to the first coordinate from the least to greatest. That is, $\delta^{1}_{1}<\delta^{2}_{1}<\ldots<\delta^{k}_{1}$, and similarly for $B$ (superscripts index the elements of $A,B$, subscripts denote coordinates). Since $A$ and $B$ are antichains it must be the case that the order on the second coordinate is opposite, i.e., in $A$, $\delta^{1}_{2}>\delta^{2}_{2},\ldots>\delta^{k}_{2}$ and similarly in $B$. Since $A$ hits the minimal value on both coordinates and $B$ hits the maximal value on both coordinates it follows that $\delta^{1}_{1}$ is the minimum value on the first coordinate in $\Theta$, whereas $\theta^{1}_{2}$ is the maximal value on the second coordinate in $\Theta$. In particular, we have $\delta^{1}_{1}\leq\theta^{1}_{1}$ and $\delta^{1}_{2}\leq\theta^{1}_{2},$ so $\delta^{1}\leq\theta^{1}$.
Next, we claim that one of the following holds
To see that the third condition in Definition 9 holds, assume towards contradiction that $A\cup B^{\prime}$ is compensatory (in $\Theta^{\prime}$). Let $\bar{\theta}^{\prime}=(\bar{\theta}_{1},\theta^{2}_{2})$ denote the maximal value in $\Theta^{\prime}$. Let $\theta^{*}\in\Theta^{\prime}$ be such that $\theta^{*}\ll\bar{\theta}^{\prime}$ and such that every state in $A\cup B^{\prime}$ is neither weakly below $\theta^{*}$ nor strictly above it. This implies that $\theta^{2}_{1}\leq\theta^{*}_{1}$ (otherwise $\theta^{2}\gg\theta^{*}$), which in turn implies that $\theta^{1}_{1}<\theta^{*}_{1}$. But we also have $\theta^{1}_{2}>\theta^{*}_{2}$, so $\theta^{1}$ is neither weakly below $\theta^{*}$ nor strictly above it. But then every state in $A\cup B$ is neither weakly below $\theta^{*}$ nor strictly above it and $\theta^{*}\ll\bar{\theta}$, so $A\cup B$ is compensatory in $\Theta$ in contradiction to $B>^{\Theta}_{AC}A$.
We have established that $B^{\prime}>^{\Theta^{\prime}}_{AC}A$. Since $|\Theta^{\prime}|<s$, by the induction hypothesis there exist distributions $P^{A},P^{B^{\prime}}$ supported on $A$ and $B^{\prime}$, respectively, such that $P^{A}\prec^{*}_{cw}P^{B^{\prime}}$ on $\Theta^{\prime}$. Recall that $\underaccent{\bar}{\theta}$ is the minimal state in $\Theta$ and $\bar{\theta}^{\prime}=(\bar{\theta}_{1},\theta^{2}_{2})$ is the maximal state in $\Theta^{\prime}$. So $P^{A}\prec^{*}_{cw}P^{B^{\prime}}$ implies that there exists $\varepsilon>0$ such that for every coordinate $i$ and for every $t\in\Theta^{\prime}_{i}$ with $t<\max(\Theta^{\prime}_{i})$ we have $F^{A}_{i}(t)>F^{B^{\prime}}_{i}(t)+2\varepsilon$. Let $D_{\theta^{1}}$ be the Dirac measure on $\theta^{1}$ and define the mixture distribution $P^{B}=(1-\varepsilon)P^{B^{\prime}}+\varepsilon D_{\theta^{1}}$. Then, for any $x\in\Theta^{\prime}_{1}=\Theta_{1}$, with $x<\bar{\theta}^{\prime}_{1}=\bar{\theta}_{1}$ we have
So, $P^{A}_{1}\prec^{*}_{st}P^{B}_{1}$ in $\Theta_{1}$.
Moving on to the second coordinate, fix any $y\in\Theta_{2}$ with $y<\bar{\theta}_{2}$. Since $\theta^{1}_{2}=\bar{\theta}_{2}$, the Dirac mass on $\theta^{1}$ assigns no probability to $\{\theta_{2}\leq y\}$, and therefore
$F^{B}_{2}(y)=(1-\varepsilon)F^{B^{\prime}}_{2}(y).$
If $F^{B^{\prime}}_{2}(y)>0$, then $(1-\varepsilon)F^{B^{\prime}}_{2}(y)<F^{B^{\prime}}_{2}(y)\leq F^{A}_{2}(y)$. If instead $F^{B^{\prime}}_{2}(y)=0$, then $F^{B}_{2}(y)=0$ while $F^{A}_{2}(y)>0$ because $P^{A}$ has full support on $A$ and $A$ hits the minimal value on the second coordinate, so $P^{A}_{2}(\underaccent{\bar}{\theta}_{2})>0$. In either case, we have $F^{B}_{2}(y)<F^{A}_{2}(y)$.
Hence, $P^{A}_{2}\prec^{*}_{st}P^{B}_{2}$ on $\Theta_{2}$ and therefore $P^{A}\prec^{*}_{cw}P^{B}$ on $\Theta$, which completes the induction argument.
Proving that $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma)$:
  1. 1.
    Recall that $\Theta_{1}=\{x_{1},\ldots,x_{m}\}$ with $x_{1}<x_{2}<\ldots<x_{m}$. The assumption that $\theta$ is not minimal on the first coordinate implies $\theta_{1}=x_{j}$ for some $1<j\leq m$. Define $\theta^{*}=(x_{j-1},\theta_{2})$. Then $\theta^{*}\ll\bar{\theta}$. We will show that for every $\theta^{\prime}\in\Gamma$ it is not the case that $\theta^{\prime}\leq\theta^{*}$ and it is not the case that $\theta^{\prime}\gg\theta^{*}$ so $\Gamma$ is compensatory in contradiction to the third condition in Theorem 2.
  2. 2.
  3. 3.
    $\min(\Gamma)\cup\max(\Gamma)$ is non-compensatory: Assume towards contradiction that there exists $\theta\in\Theta$, with $\theta\ll\bar{\theta}$, such that for every $\theta^{\prime}\in\min(\Gamma)\cup\max(\Gamma)$ it is not the case that $\theta^{\prime}\leq\theta$ and it is not the case that $\theta^{\prime}\gg\theta$. Then there does not exist $\theta^{\prime}\in\Gamma$ with $\theta^{\prime}\gg\theta$ since that would imply the existence of $\tilde{\theta}\in\max(\Gamma)$ such that $\tilde{\theta}\geq\theta^{\prime}\gg\theta$, a contradiction. Similarly, there does not exist $\theta^{\prime}\in\Gamma$ with $\theta^{\prime}\leq\theta$ as that would imply the existence of $\tilde{\theta}\in\min(\Gamma)$ such that $\tilde{\theta}\leq\theta^{\prime}\leq\theta$. So $\Gamma$ is compensatory, a contradiction.
Proving that $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma^{C})$:
  1. 1.
    $\min(\Gamma)$ and $\max(\Gamma^{C})$ are disjoint: immediate.
  2. 2.
  3. 3.
    $\min(\Gamma)\cup\max(\Gamma^{C})$ is non-compensatory: Assume towards contradiction that there exists $\theta\in\Theta$, with $\theta\ll\bar{\theta}$, such that for every $\theta^{\prime}\in\min(\Gamma)\cup\max(\Gamma^{C})$ it is not the case that $\theta^{\prime}\leq\theta$ and it is not the case that $\theta^{\prime}\gg\theta$. Then $\Gamma$ includes all states strictly above $\theta$, because if $\theta^{\prime}\gg\theta$ and $\theta^{\prime}\in\Gamma^{C}$ then there exists some $\tilde{\theta}\in\max(\Gamma^{C})$ such that $\tilde{\theta}\geq\theta^{\prime}\gg\theta$, a contradiction. Similarly, $\Gamma^{C}$ includes all states weakly below $\theta$, as otherwise there exists $\tilde{\theta}\in\min(\Gamma)$ with $\tilde{\theta}\leq\theta$. Thus, $\Gamma$ is biased upward, in contradiction to the second condition in Theorem 2.
Proving that $\min(\Gamma^{C})<^{\Theta}_{AC}\max(\Gamma)$: The arguments for the first two conditions of Definition 9 are exactly as those in the proof that $\min(\Gamma)<^{\Theta}_{AC}\max(\Gamma^{C})$. For the third condition, assume towards contradiction that there exists $\theta\in\Theta$, with $\theta\gg\underaccent{\bar}{\theta}$, such that for every $\theta^{\prime}\in\min(\Gamma^{C})\cup\max(\Gamma)$ it is not the case that $\theta^{\prime}\geq\theta$ and it is not the case that $\theta^{\prime}\ll\theta$. [^30] Here we are using the equivalent definition for a set to be compensatory that is stated in Footnote 15. Then, $\Gamma$ includes all states strictly below $\theta$, because if there exists $\theta^{\prime}\in\Gamma^{C}$ with $\theta^{\prime}\ll\theta$ then there exists $\tilde{\theta}\in\min(\Gamma^{C})$ such that $\tilde{\theta}\leq\theta^{\prime}\ll\theta$, a contradiction. Similarly, $\Gamma^{C}$ includes all states weakly above $\theta$. Thus, $\Gamma$ is biased downward in contradiction to the second condition in Theorem 2.
Appendix B Probability of Polarization